Circular No. 6243

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SATELLITES OF SATURN P. D. Nicholson and C. A. McGhee, Cornell University; M. R. Showal-

ter and L. Dones, Ames Research Center, NASA; R. G. French, Wellesley College; S. M. Larson, Lunar and Planetary Laboratory; J. J. Lissauer, State University of New York, Stony Brook; P. O. Seitzer, Unversity of Michigan; B. Sicardy, Université de Paris; and G. E. Danielson, California Institute of Technology, report: “During a 12-hr period spanning the Aug. 10 Saturnian ring-plane crossing, we obtained a total of 48 images with the HST Wide Field Planetary Camera 2 (+ 890-nm methane-band filter). In addition to Saturn X (Janus), XI (Epimetheus), and XVII (Pandora), the images reveal the presence of at least four other objects orbiting in the vicinity of the F ring. We confirm that S/1995 S 2 lags behind the predicted position of Saturn XVI (Prometheus) by 19^{o}, comparable to the 21^{o} found during the May 22 crossing (IAUC 6192), thus confirming the probable identification of S/1995 S 2 with Prometheus (IAUC 6192, 6196), though it appears to be 0.5 mag fainter than expected. For each of the other three objects we give below the orbital radius, mean daily motion, longitude at epoch 1995 Aug. 10.5 TT (at Saturn) measured from the ascending node of Saturn’s equator on the J2000.0 earth equator, equivalent radius, number of observations fitted, and approximate V magnitude at opposition: S/1995 S 5, 140 060 ± 400 km, 583^{o}.0 ± 2^{o}.5, 130^{o}.0 ± 0^{o}.5, 26 km, 15, 17.1; S/1995 S 6, 139 910 ± 200 km, 584^{o}.0 ± 1^{o}.2, 246^{o}.0 ± 0^{o}.1, 19 km, 9, 17.8; S/1995 S 7, 139 440 ± 250 km, 586^{o}.9 ± 1^{o}.6, 324^{o}.7 ± 0^{o}.3, 18 km, 10, 17.9. These fits assume circular orbits but include the effects of Saturn’s J_{2} and J_{4}. Maximum residuals are 0′′.16 (or 1.6 pixels in the Wide Field Camera), corresponding to 1000 km or less. Radii are calculated from the objects’ integrated fluxes, assuming circular cross-sections and albedos similar to those of Saturn X and XI. The orbits of S/1995 S 5 and S/1995 S 6 are consistent with that of the F ring (a = 140 200 km, n = 582^{o}.3 day^{−1}), and we suspect that these objects are clumps or arcs within this ring rather than undiscovered satellites, based on their brightness and the completeness limit of Voyager imaging searches (R ≈ 10 km). S/1995 S 5 shows appreciable brightness variations with orbital phase, also suggesting an arc-like structure. The orbit of S/1995 S 7 is indistinguishable from that of Prometheus = S/1995 S 2, with the new object trailing the latter by 15^{o}. The above orbital radii are substantially different from those of S/1995 S 1 and S/1995 S 3 (IAUC 6192) and their mean motions too imprecise at present to permit unambiguous linkage with these objects.”

1995 October 4 Daniel W. E. Green